OpenStudy (anonymous):

Solve 2x + 5 = 27 A. 8.5 B. 64 C. 16 D. 11

OpenStudy (anonymous):

I thought it was 20 but theres no 20 was all like

OpenStudy (anonymous):

2x+5-5=27-5

OpenStudy (ageta):

2x + 5 = 27 -5 -5 2x=22 x=11

OpenStudy (ageta):

x=11

OpenStudy (anonymous):

oh

sammixboo (sammixboo):

\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm W}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm L}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm C}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm M}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\) \(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\) \(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm P}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm N}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm S}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm U}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm D}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm Y}}}\) To solve this we must first get x alone, so subtract 5 on both sides of the equation 2x + 5 - 5 = 27 - 5 2x = 27 - 5 Can you tell me what is 27 - 5?

OpenStudy (anonymous):

ya its 22

sammixboo (sammixboo):

Good! So now we have 2x = 22 To finally make x alone we must divide 2 on both sides of the equation 2x/2 = 22/2 x = 22/2 So what is 22/2?

OpenStudy (anonymous):

ll

sammixboo (sammixboo):

Yes, and that is your answer :)

OpenStudy (anonymous):

ohh

sammixboo (sammixboo):

Good job :P

OpenStudy (anonymous):

thanx

OpenStudy (ageta):

i just taught him

OpenStudy (ageta):

your welcome

OpenStudy (anonymous):

her

sammixboo (sammixboo):

No problem! If you need anymore help you can tag me in your question @sammixboo

OpenStudy (ageta):

i just answer it.

OpenStudy (anonymous):

okay thanx

OpenStudy (ageta):

you welcome

sammixboo (sammixboo):

It is always best to explain each step to solving the Asker's question, so you can show the Asker how to do future problems like this one :)

OpenStudy (anonymous):

What is the value of x? x/5+6 =36 x = ____ by the way the x/5 is a fraction

OpenStudy (anonymous):

i got 150 as my answer

OpenStudy (anonymous):

@ageta